OpenStudy (anonymous):

the circumference of a circle exceed its diameter by 90cm .Find the area of the circle @Werewolfprincess

OpenStudy (anonymous):

eek math!

OpenStudy (some.random.cool.kid):

oooo math :}

OpenStudy (cwrw238):

circumference = pi * d where d = diameter pi * d - d = 90 solve for d radius r = d /2 area = pi r^2

OpenStudy (anonymous):

?

OpenStudy (cwrw238):

first find the diameter of the circle by solving the equation pi*d - d = 90 d( pi - 1) = 90 d = 90 / (pi - 1) work this out on ur calculator then divide by 2 to find the radius of the circle (r) then find the area using the formula pi * r^2

OpenStudy (anonymous):

we are not allowed to use a calci

OpenStudy (cwrw238):

ok - use 3.14 for pi so diameter d = 90 / (3.14 - 1) = 90 / 2.14

OpenStudy (cwrw238):

you'll have to use long division to find d